# Kinematics of Circumgalactic OĀ VIĀ Gas and Disk Rotation of Star-forming Galaxies

### v1

@@ 3.3 Comparison Between OĀ VIĀ and Low-ionization-state Gas Kinematics
Not only the SiĀ II- and SiĀ III-matched OĀ VIĀ components rarely counter-rotate, but they also preferentially corotate with the disk instead of being at the galaxy systemic velocity. Specifically, 9/10 (90%) of the SiĀ II-matched OĀ VIĀ components corotate with the disk in contrast to only 7/19 (37%) of the OĀ VIĀ components without a SiĀ IIĀ match. The contrast becomes less prominent for SiĀ III. But still, a higher fraction (10/14; 71%) of the SiĀ III-matched OĀ VIĀ components corotate with the disk compared to that of the no-SiĀ III-matched OĀ VIĀ components (6/15; 40%). Our results show that the OĀ VIĀ velocity components with and without low-ion matches have different kinematics properties and possibly trace gas that resides at different 3D distances from the host galaxies. Because OĀ VIĀ velocity components with matched low-ion are more likely to corotate with the disk, a plausible scenario is that the gas resides on an extended disk plane. We further discuss the implications of our results in SectionĀ 4.
@@ 4 Discussion
In SectionĀ 3, we have presented the kinematics of the OĀ VIĀ gas and compared that with galaxy disk rotation and the kinematics of low-ionization-state gas traced by SiĀ IIĀ and SiĀ III. While individual velocity components of the OĀ VIĀ gas do not seem to correlate with the galaxy disk rotation, the column-density-weighted velocity of OĀ VIĀ suggests that the bulk of OĀ VIĀ gas rarely counter-rotate with the disk (FigureĀ 5). We have also matched velocity components of OĀ VIĀ with SiĀ IIĀ and SiĀ IIIĀ by minimizing the differences of their velocity centroids. This analysis has shown that the SiĀ II- or SiĀ III-matched OĀ VIĀ velocity components more frequently corotate with the disk compared to velocity components that cannot be matched with the low ions (FigureĀ 9). In this Section, we discuss the implications and interpretation of these results.
@@ 4.1 Location of the Multiphase Gas Traced by OĀ VI, SiĀ II, and SiĀ III
Among the 12 sightlines with OĀ VIĀ detection and spectral coverage of SiĀ IIĀ and SiĀ III, OĀ VIĀ is detected in sightlines out to larger impact parameters compared to the low ions (FiguresĀ 5, 6, and 8). This result is commonly found in previous work. While low ions are only detected in the inner CGM out to kpc around star-forming galaxies, OĀ VIĀ is detected out to impact parameters of several hundred kpcs (see SectionĀ 1). This implies that OĀ VIĀ is more spatially extended compared to the low-ionization-state gas that resides in the inner CGM.
@@
In general, quasar absorption-line analysis only produces line-of-sight measurements and do not reveal the location of the absorbing gas in 3D space. However, our kinematics comparison between OĀ VIĀ and low ions provides the clues for distinguishing the location of the detected gas. FigureĀ 9 shows that SiĀ II- and SiĀ III-matched OĀ VIĀ velocity components more frequently corotate with the disk compared to those components without low-ion matches, and most of the matched components are found in sightlines with impact parameters below 100 kpc. This may not seem surprising, because low ions are often detected in the inner CGM and corotate with the disk, suggesting that the low ions trace the gas residing near the extended disk plane of the galaxy (e.g., Kacprzak etĀ al., 2011; Ho etĀ al., 2017). But with the subset of OĀ VIĀ components having velocities that matched with the low ions and also corotating with the disk, this suggests that these low-ion-matched OĀ VIĀ components trace the gas that is co-spatial with the low ions, i.e., near the extended disk plane in the inner CGM. On the other hand, the unmatched OĀ VIĀ velocity components correspond to gas at the outer CGM.
@@
The interpretation of the location of the absorbing gas explains two results. First, it explains why the SiĀ II- and SiĀ III-matched OĀ VIĀ velocity components are more likely to be corotating with the disk compared to those OĀ VIĀ components without a low-ion match. It also explains the complicated line structures detected in sightlines with small impact parameters (FigureĀ 8), because these sightlines intersect both the inner and outer CGM with different OĀ VIĀ kinematics. Therefore, our results imply that while the outer OĀ VIĀ CGM tends to move randomly, the inner OĀ VIĀ CGM carries angular momentum and corotates with the galaxy disk similar to the low-ionization-state gas.
@@
Furthermore, the OĀ VIĀ velocity components with SiĀ IIĀ matches tend to have larger Doppler parameters compared to those without SiĀ IIĀ matches. FigureĀ 10 shows the distributions of the Doppler parameter and column density for the two OĀ VIĀ populations. Performing the Kolmogorov-Smirnov (KS) test on the Doppler parameter rules out the null hypothesis at 96% that the OĀ VIĀ components with and without SiĀ IIĀ matches are drawn from the same population distribution. However, the null hypothesis cannot be rejected for the column density populations. On the other hand, when we repeat the same analysis for the OĀ VIĀ components with and without SiĀ IIIĀ matches, the null hypothsis cannot be ruled out for both measured quantities. In fact, since SiĀ II-matched OĀ VIĀ components are more likely to be corotating with the disk compared to those without SiĀ IIĀ matches, the larger Doppler parameter of the former plausibly suggests that the rotation motion (in addition to turbulence) has broadened the absorption lines. Because the net CGM angular momentum remains roughly constant with radius in 3D space, the rotation velocity is expected to decrease towards the outer CGM. With SiĀ IIIĀ generally being more spatially extended compared to the lower-ionization SiĀ IIĀ counterpart, we expect the OĀ VIĀ components with SiĀ IIIĀ matches to trace the gas out to larger 3D radii and have lower rotation velocities on average compared to those with SiĀ IIĀ matches. As a result, the rotation motion has produced a measurable increase to the absorption-line width of the OĀ VIĀ velocity components with SiĀ IIĀ matches but not those with SiĀ IIIĀ matches, which also explains why the latter have comparable Doppler parameters as the no-SiĀ III-matched OĀ VIĀ components.
@@ 4.2 Angular Momentum of the CGM and Radial Infall
While the absorption-line measurements do not reveal where the gas is located along the line of sight, we have argued that some of the OĀ VIĀ velocity components are tracing gas near the extended disk plane of galaxies (SectionĀ 4.1). But does the corotating gas have sufficient angular momentum to stay on circular orbits? If not, the gas could be infalling to the galaxy, a plausible scenario that explains the corotating low-ionization-state gas at sub-centrifugal velocities (e.g., Ho etĀ al., 2017; Ho & Martin, 2020).
We explore this radial inflow hypothesis by assuming all the non-counter-rotating gas being on the extended disk plane. Then, we calculate the implied galactocentric radii and the circular velocities for individual SiĀ II, SiĀ III, and OĀ VIĀ velocity components. We compare these circular velocities to that required for the gas to stay on circular orbits, assuming the galaxy disks maintain the flat, arctangent rotation curve out to the galactocentric radii. The histograms in FigureĀ 11 show the fractions of SiĀ II, SiĀ III, and OĀ VIĀ velocity components with circular velocities too low to be on circular orbits, i.e., the gas lacks sufficient angular momentum and is infalling (filled histograms)131313 Velocity components that are consistent with being at the galaxy systemic velocities are considered as lacking sufficient angular momentum to stay on circular orbits. . The fraction of infalling velocity components clearly increases with decreasing ionization potential, for which only about 45% for OĀ VIĀ velocity components are tracing infalling gas, and this fraction increases to 80% for SiĀ II. On the other hand, about 15% of the velocity components of each ion have excess angular momentum to be on the circular orbits (hatched). This gas is not expected to be accreting onto the disk.
With the infall interpretation, we can also determine the radii where the infalling gas joins the disk. Assuming the specific angular momentum of the gas remains unchanged, i.e., stays constant, we calculate the radius where the gas has the same specific angular momentum as the extended disk (also known as the circularization radius). For SiĀ II, SiĀ III, and OĀ VIĀ gas, the calculation produces typical radii of under 50 kpc, which is comparable to the radii of extraplanar HĀ IĀ gas of nearby disk galaxies from 21-cm emission observations (e.g., Marasco etĀ al., 2019).
@@
Therefore, we repeat the analysis by comparing the circular velocities required for the gas to stay on circular orbits to the column-density-weighted velocities of different ions in individual sightlines. This enables us to consider the average angular momentum of the gas intersected along the sightlines instead of the gas clouds associated with individual velocity components. The result is shown in FigureĀ 12. First of all, the fraction of counter-rotating OĀ VIĀ gas is only about 10%, which is four times lower than that calculated using individual velocity components (%). This reiterates the result from FigureĀ 5 that net counter-rotating OĀ VIĀ absorbers are rare. Also, the infalling gas fractions of SiĀ II, SiĀ III, and OĀ VIĀ gas all reach %, which are higher than the calculations with individual velocity components, especially for OĀ VIĀ (45%). Because OĀ VIĀ traces gas out to larger 3D radii compared to the low ions, and the expected CGM rotation velocity decreases with increasing radius due to the net CGM angular momentum being roughly radius-independent (also see SectionĀ 4.1), the OĀ VIĀ corotation and/or infall signatures could be easily masked by the turbulent motion. Consequently, the corotation and/or infall of OĀ VIĀ gas is more likely to be observed only when we consider the average angular momentum using the column-density-weighted velocities in individual sightlines instead of individual velocity components.
4.3 Comparison with previous work on OĀ VIĀ kinematics and disk rotation
@@
Unlike the numerous literatures that analyzed the relationship between low-ionization-state MgĀ IIĀ gas and disk rotation, very few work focused on this comparison with OĀ VIĀ (see SectionĀ 1). Using 10 major-axis sightlines with azimuthal angle below 25ā, Kacprzak etĀ al. (2019) showed that both the individual velocity components and weighted velocities per sightline of the OĀ VIĀ gas does not correlate with the disk rotation kinematics (their FigureĀ 4). They also demonstrated that a simple monolithic rotating halo model failed to explain the full velocity spread of the OĀ VIĀ absorption. In contrast, when they ran AMR cosmological simulations, they clearly found infalling OĀ VIĀ filaments that rotate in the same direction as the galaxy disks. The inflowing OĀ VIĀ gas decelerates when it approaches the galaxies and becomes rotationally-dominated. However, the inflowing/rotating OĀ VIĀ signatures were washed out by other velocity structures in the CGM when they computed integrated velocities along lines of sight through the entire simulated galaxy halos. Therefore, inspired by their simulation analysis, they concluded that while corotating or accreting OĀ VIĀ gas could still exist in their observed sightlines, the signature was either weak or was masked by other kinematics ongoing within the CGM.
@@
More recently, Kacprzak etĀ al. (2025) conducted the COS-EDGES survey and analyzed the multiphase CGM of nine edge-on disk galaxies (ā) probed by major-axis sightlines at impact parameters below 0.3. By comparing the stacked absorption-line profiles of different ions (e.g., MgĀ I, MgĀ II, and OĀ VI) in the small () and large () impact parameter bins, they showed that while the Doppler sign of the weighted velocities are consistent with disk rotation, the velocity magnitude decreases with increasing impact parameters. More importantly, the amount of absorption that corotates with the disk declines with both increasing ionization potential and sightline impact parameters; in contrast to over 90% of the MgĀ IIĀ absorption that corotates with the disk, only 80% (60%) of the OĀ VIĀ absorption corotates at small (large) impact parameters. Hence, they concluded that the CGM has a radially-dependent kinematic structure. Within 0.2, the CGM hosts cool and dynamically broad gas that tightly coupled with disk rotation, whereas beyond 0.2, especially for OĀ VI, the CGM becomes less kinematically correlated with disk rotation due to the presence of corotating, lagging, and volume-filling components.
@@
In general, our analysis agrees with both Kacprzak etĀ al. (2019) and Kacprzak etĀ al. (2025) that the corotation signature of OĀ VIĀ gas is challenging to observe due to kinematics components other than rotation. In particular, while individual OĀ VIĀ velocity components do not seem to correlate with disk rotation, the column-density-weighted velocities suggest that the bulk of OĀ VIĀ gas rarely counter-rotates (FigureĀ 5). We have attributed this ādiscrepancyā to the turbulent motion of the CGM, and averaging is required to reveal the net OĀ VIĀ CGM angular momentum while searching for the corotation signature (SectionĀ 4.2). Furthermore, we have also compared the OĀ VIĀ kinematics to that of the low ions. Not only do low ions more often corotate with the disk compared to OĀ VIĀ (FigureĀ 6), but the OĀ VIĀ velocity components with low-ion matches are more likely to corotate compared to those without low-ion matches (FigureĀ 8). These results suggest that low ions are better indicators for identifying the corotating gas, which also matches the conclusion from Nateghi etĀ al. (2024) and Kacprzak etĀ al. (2025) that the lower the ionization potential, the more likely the ion is kinematically consistent with the rotation of the galaxy disks and supports the accretion scenario.

### v2

@@ 3.3 Comparison Between OĀ VI and Low-ionization-state Gas Kinematics
Not only the SiĀ II- and SiĀ III-matched OĀ VI components rarely counter-rotate, but they are also more likely to corotate with the disk compared to the unmatched OĀ VI velocity components. Specifically, 9/10 (90%) of the SiĀ II-matched OĀ VI components corotate with the disk in contrast to only 7/19 (37%) of the OĀ VI components without a SiĀ II match. The contrast becomes less prominent for SiĀ III. But still, a higher fraction (10/14; 71%) of the SiĀ III-matched OĀ VI components corotate with the disk compared to that of the no-SiĀ III-matched OĀ VI components (6/15; 40%). Our results show that the OĀ VI velocity components with and without low-ion matches have different kinematics properties and possibly trace gas that resides at different 3D distances from the host galaxies. Because OĀ VI velocity components with matched low-ion are more likely to corotate with the disk, a plausible scenario is that the gas resides on an extended disk plane. We further discuss the implications of our results in SectionĀ 4.
@@ 4 Discussion
In SectionĀ 3, we have presented the kinematics of the OĀ VI gas and compared that with galaxy disk rotation and the kinematics of low-ionization-state gas traced by SiĀ II and SiĀ III. While individual velocity components of the OĀ VI gas do not seem to correlate with the galaxy disk rotation, the column-density-weighted velocity of OĀ VI suggests that the bulk of OĀ VI gas rarely counter-rotate with the disk (FigureĀ 5). We have also matched velocity components of OĀ VI with SiĀ II and SiĀ III by minimizing the differences of their velocity centroids. This analysis has shown that the SiĀ II- or SiĀ III-matched OĀ VI velocity components more frequently corotate with the disk compared to velocity components that cannot be matched with the low ions (FigureĀ 9). In this Section, we discuss the implications and interpretation of these results.
@@ 4.1 Location of the Multiphase Gas Traced by OĀ VI, SiĀ II, and SiĀ III
Among the 12 sightlines with OĀ VI detection and spectral coverage of SiĀ II and SiĀ III, OĀ VI is detected in sightlines out to larger impact parameters compared to the low ions (FiguresĀ 5, 6, and 8). This result is commonly found in previous work. While low ions are only detected in the inner CGM out to kpc around star-forming galaxies, OĀ VI is detected out to impact parameters of several hundred kpcs (see SectionĀ 1). This implies that OĀ VI is more spatially extended compared to the low-ionization-state gas that resides in the inner CGM.
@@
In general, quasar absorption-line analysis only produces line-of-sight measurements and do not reveal the location of the absorbing gas in 3D space. However, our kinematics comparison between OĀ VI and low ions provides the clues for distinguishing the location of the detected gas. FigureĀ 9 shows that SiĀ II- and SiĀ III-matched OĀ VI velocity components more frequently corotate with the disk compared to those components without low-ion matches, and most of the matched components are found in sightlines with impact parameters below 100 kpc. This may not seem surprising, because low ions are often detected in the inner CGM and corotate with the disk, suggesting that the low ions trace the gas residing near the extended disk plane of the galaxy (e.g., Kacprzak etĀ al., 2011; Ho etĀ al., 2017). But with the subset of OĀ VI components having velocities that matched with the low ions and also corotating with the disk, this suggests that these low-ion-matched OĀ VI components trace the gas that is co-spatial with the low ions (also see Werk etĀ al., 2016), i.e., near the extended disk plane in the inner CGM. On the other hand, the unmatched OĀ VI velocity components correspond to gas at the outer CGM. In fact, Werk etĀ al. (2016) concluded that the unmatched OĀ VI components trace the volume-filling hot corona at the halo virial temperature, and our results do not contradict with that description.
@@
The interpretation of the location of the absorbing gas explains two results. First, it explains why the SiĀ II- and SiĀ III-matched OĀ VI velocity components are more likely to be corotating with the disk compared to those OĀ VI components without a low-ion match. It also explains the complicated line structures detected in sightlines with small impact parameters (FigureĀ 8), because these sightlines intersect both the inner and outer CGM with different OĀ VI kinematics. This is similar to the result from multi-sightline probes of the Andromeda Galaxy that demonstrated the decreasing kinematic complexity of the CGM with increasing radii (Lehner etĀ al., 2025). Therefore, our results imply that while the outer OĀ VI CGM tends to move randomly, the inner OĀ VI CGM carries angular momentum and corotates with the galaxy disk similar to the low-ionization-state gas.
@@
Furthermore, the OĀ VI velocity components with SiĀ II matches tend to have larger Doppler parameters compared to those without SiĀ II matches. FigureĀ 10 shows the distributions of the Doppler parameter and column density for the two OĀ VI populations. Performing the Kolmogorov-Smirnov (KS) test on the Doppler parameter rules out the null hypothesis at 96% that the OĀ VI components with and without SiĀ II matches are drawn from the same population distribution. However, the null hypothesis cannot be rejected for the column density populations. On the other hand, when we repeat the same analysis for the OĀ VI components with and without SiĀ III matches, the null hypothsis cannot be ruled out for both measured quantities. In fact, since SiĀ II-matched OĀ VI components are more likely to be corotating with the disk compared to those without SiĀ II matches, the larger Doppler parameter of the former plausibly suggests that the rotation motion (in addition to turbulence) has broadened the absorption lines. Because the net CGM angular momentum remains roughly constant with radius in 3D space, the rotation velocity is expected to decrease towards the outer CGM. With SiĀ III generally being more spatially extended compared to the lower-ionization SiĀ II counterpart, we expect the OĀ VI components with SiĀ III matches to trace the gas out to larger 3D radii and have lower rotation velocities on average compared to those with SiĀ II matches. As a result, the rotation motion has produced a measurable increase to the absorption-line width of the OĀ VI velocity components with SiĀ II matches but not those with SiĀ III matches, which also explains why the latter have comparable Doppler parameters as the no-SiĀ III-matched OĀ VI components.
@@ 4.2 Angular Momentum of the CGM and Radial Infall
While the absorption-line measurements do not reveal where the gas is located along the line of sight, we have argued that some of the OĀ VI velocity components are tracing gas near the extended disk plane of galaxies (SectionĀ 4.1). But does the corotating gas have sufficient angular momentum to stay on circular orbits? If not, the gas could be infalling to the galaxy, a plausible scenario that explains the corotating low-ionization-state gas at sub-centrifugal velocities (e.g., Ho etĀ al., 2017; Ho & Martin, 2020).
We explore this radial inflow hypothesis by assuming all the non-counter-rotating gas being on the extended disk plane. Then, we calculate the implied galactocentric radii and the circular velocities for individual SiĀ II, SiĀ III, and OĀ VI velocity components. We compare these circular velocities to that required for the gas to stay on circular orbits, assuming the galaxy disks maintain the flat, arctangent rotation curve out to the galactocentric radii. The histograms in FigureĀ 11 show the fractions of SiĀ II, SiĀ III, and OĀ VI velocity components with circular velocities too low to be on circular orbits, i.e., the gas lacks sufficient angular momentum and is infalling (filled histograms)131313Velocity components that are consistent with being at the galaxy systemic velocities are considered as lacking sufficient angular momentum to stay on circular orbits. . The fraction of infalling velocity components clearly increases with decreasing ionization potential, for which only about 45% for OĀ VI velocity components are tracing infalling gas, and this fraction increases to 80% for SiĀ II. On the other hand, about 15% of the velocity components of each ion have excess angular momentum to be on the circular orbits (hatched). This gas is not expected to be accreting onto the disk.
With the infall interpretation, we can also determine the radii where the infalling gas joins the disk. Assuming the specific angular momentum of the gas remains unchanged, i.e., stays constant, we calculate the radius where the gas has the same specific angular momentum as the extended disk (also known as the circularization radius). For SiĀ II, SiĀ III, and OĀ VI gas, the calculation produces typical radii of under 50 kpc, which is comparable to the radii of extraplanar HĀ I gas of nearby disk galaxies from 21-cm emission observations (e.g., Marasco etĀ al., 2019).
@@
Therefore, we repeat the analysis by comparing the circular velocities required for the gas to stay on circular orbits to the column-density-weighted velocities of different ions in individual sightlines. This enables us to consider the average angular momentum of the gas intersected along the sightlines instead of the gas clouds associated with individual velocity components. The result is shown in FigureĀ 12. First of all, the fraction of counter-rotating OĀ VI gas is only about 10%, which is four times lower than that calculated using individual velocity components (%). This reiterates the result from FigureĀ 5 that net counter-rotating OĀ VI absorbers are rare. Also, the infalling gas fractions of SiĀ II, SiĀ III, and OĀ VI gas all reach %, which are higher than the calculations with individual velocity components, especially for OĀ VI (45%). Because OĀ VI traces gas out to larger 3D radii compared to the low ions, and the expected CGM rotation velocity decreases with increasing radius due to the net CGM angular momentum being roughly radius-independent (also see SectionĀ 4.1), the OĀ VI corotation and/or infall signatures could be easily masked by the turbulent motion. Consequently, the corotation and/or infall of OĀ VI gas is more likely to be observed only when we consider the average angular momentum using the column-density-weighted velocities in individual sightlines instead of individual velocity components.
4.3 Comparison with previous work on OĀ VI kinematics and disk rotation
@@
Unlike the numerous literatures that analyzed the relationship between low-ionization-state MgĀ II gas and disk rotation, very few work focused on this comparison with OĀ VI (see SectionĀ 1). Using 10 major-axis sightlines with azimuthal angle below 25ā, Kacprzak etĀ al. (2019) showed that both the individual velocity components and weighted velocities per sightline of the OĀ VI gas does not correlate with the disk rotation kinematics (their FigureĀ 4). They also demonstrated that a simple monolithic rotating halo model failed to explain the full velocity spread of the OĀ VI absorption. In contrast, when they ran AMR cosmological simulations, they clearly found infalling OĀ VI filaments that rotate in the same direction as the galaxy disks. The inflowing OĀ VI gas decelerates when it approaches the galaxies and becomes rotationally-dominated. However, the inflowing/rotating OĀ VI signatures were washed out by other velocity structures in the CGM when they computed integrated velocities along lines of sight through the entire simulated galaxy halos. Therefore, inspired by their simulation analysis, they concluded that while corotating or accreting OĀ VI gas could still exist in their observed sightlines, the signature was either weak or was masked by other kinematics ongoing within the CGM.
@@
More recently, Kacprzak etĀ al. (2025) conducted the COS-EDGES survey and analyzed the multiphase CGM of nine edge-on disk galaxies (ā) probed by major-axis sightlines at impact parameters below 0.3. By comparing the stacked absorption-line profiles of different ions (e.g., MgĀ I, MgĀ II, and OĀ VI) in the small () and large () impact parameter bins, they showed that while the Doppler sign of the weighted velocities are consistent with disk rotation, the velocity magnitude decreases with increasing impact parameters. More importantly, the amount of absorption that corotates with the disk declines with both increasing ionization potential and sightline impact parameters; in contrast to over 90% of the MgĀ II absorption that corotates with the disk, only 80% (60%) of the OĀ VI absorption corotates at small (large) impact parameters. Hence, they concluded that the CGM has a radially-dependent kinematic structure. Within 0.2, the CGM hosts cool and dynamically broad gas that tightly coupled with disk rotation, whereas beyond 0.2, especially for OĀ VI, the CGM becomes less kinematically correlated with disk rotation due to the presence of corotating, lagging, and volume-filling components.
@@
In general, our analysis agrees with both Kacprzak etĀ al. (2019) and Kacprzak etĀ al. (2025) that the corotation signature of OĀ VI gas is challenging to observe due to kinematics components other than rotation. In particular, while individual OĀ VI velocity components do not seem to correlate with disk rotation, the column-density-weighted velocities suggest that the bulk of OĀ VI gas rarely counter-rotates (FigureĀ 5). We have attributed this ādiscrepancyā to the turbulent motion of the CGM, and averaging is required to reveal the net OĀ VI CGM angular momentum while searching for the corotation signature (SectionĀ 4.2). Furthermore, we have also compared the OĀ VI kinematics to that of the low ions. Not only do low ions more often corotate with the disk compared to OĀ VI (FigureĀ 6), but the OĀ VI velocity components with low-ion matches are more likely to corotate compared to those without low-ion matches (FigureĀ 8). These results suggest that low ions are better indicators for identifying the corotating gas, which also matches the conclusion from Nateghi etĀ al. (2024) and Kacprzak etĀ al. (2025) that the lower the ionization potential, the more likely the ion is kinematically consistent with the rotation of the galaxy disks and supports the accretion scenario.